I need help with dosage homework. Thank you!

At 0800 the client has a continuous heparin infusion running at 12 mL per hr. According to the hospital's heparin protocol, the heparin infusion should be increased by 100 units per hour each hour. The bag hanging contains heparin 40,000 units in D5W 250 mL. Calculate the new milliliter per hour flow rate for this heparin infusion. (Round to tenths).
Answer #1

40000 units in 250 ml.

100 units in 1/40000 * 250 = 1/160 ml= 0.006 ml

At 8 ,the flow rate was 12 ml/hour.

Every hour flow rate increases by 0.006 ml/hour.

The time for which the new flow rate has to be calculated is not mentioned. So it is not possible to calculate the new flow rate.